Find the masa of an object whose rest mass is 1000g when it is travelling at 90% of the speed of light?


1
Expert's answer
2021-06-29T02:54:03-0400

The mass of an object

"m=\frac{m_0}{\sqrt{1-\frac{v^2}{c^2}}}=\frac{1000\:\rm g}{\sqrt{1-0.9^2}}=2294\:\rm g"